# Enhancing Customer Contact Efficiency with Graph Neural Networks in Credit Card Fraud Detection Workflow

### v1

@@ 1 Introduction
1.1 Our fraud detection and processing problem
@@ 1.2 Challenge
2 Related work
2.1 Bayesian Network Model
Bayesian networks excel at handling uncertainty and predicting rare events [6], both critical challenges in fraud detection. By utilizing probabilistic reasoning, they can evaluate the likelihood of fraud based on available evidence, even when data is incomplete or missing. Additionally, Bayesian networks are particularly effective in identifying rare fraud cases by incorporating prior knowledge and reasoning under uncertainty, enabling accurate predictions even when the available data is sparse or limited [3].
@@ 2.1 Bayesian Network Model
2.2 Recurrent Neural Networks
@@ 3 Methods
3.1 Dataset
The IBM credit card transaction dataset [2] is a publicly available resource designed for fraud detection research. Each record contains attributes such as transaction amount, card type, and location, along with a label indicating whether the transaction was fraudulent, the details shows at Tab. 1. This provides sufficient feature information in each row to train and evaluate models effectively. Notably, the dataset is entirely synthetic and does not correspond to real individuals or financial institutions. It comprises 24 million unique transactions, 6,000 merchants, 100,000 cards, and 30,000 fraudulent cases, accounting for 0.1% of the total transactions. Using this dataset to feed into our GNN model, we would work with approximately 16,000 nodes and 100,000 edges, representing the relationships and interactions within the transaction data.
@@ 3.1 Dataset
Table 1: Dataset Attributes and types
3.2 Architecture
@@
In this GNN, each node is characterized by two types of attributes: a concatenated string combining the card number and user name, and the merchant name represented as a string. For the edges, we primarily incorporate the attributes of each transaction as edge features 3. This approach ensures that all relevant information is effectively captured within the graph.
@@
In our model, we use RGCN (Relational Graph Convolutional Network) as the core component. RGCN is a type of GNN specifically designed for handling heterogeneous graphs, which have different types of nodes and edges. It is capable of learning on multi-relational graphs. In fraud detection scenarios, the data is often heterogeneous. Compared to traditional GCN, RGCN can model different types of relationships more accurately, improving predictive performance.
@@ 4 Experiments
The loss curve shown as figure 5. The training and testing loss curves in the figure show a sharp decline early on, indicating that the model is learning quickly. Initially, both the training and testing losses are relatively high, but they drop rapidly during the first few epochs. After around epoch 30, the losses stabilize, and both curves approach near-zero values. The training loss (blue curve) decreases more sharply and levels off, suggesting that the model is effectively learning the patterns in the data. The testing loss (orange curve) follows a similar trend but experiences more minor fluctuations, which is typical due to the nature of the testing set. Overall, the steady decrease in loss values across both training and testing phases demonstrates that the model is converging well within the first 200 epochs, even though the testing loss shows slight variability. The final accuracy can achieve 0.9988%, which is highly approach 100%.

### v2

@@ I Introduction
I-A Fraud Detection and Processing Problem
@@ I-B Challenge
Recent progress in data-driven modeling and learning-based systems has significantly enhanced our ability to address complex problems across domains. For example, nearest neighbor methods have seen renewed interest, both in density estimation [16] and in reinforcement learning settings with minimax guarantees [17]. In the context of retrieval-augmented generation, RAG-Instruct [8] explores how diverse prompts can guide large language models (LLMs) more effectively. Similarly, BlendSQL [5] offers a unified structured language to support hybrid question answering over structured and unstructured data, addressing the limitations of standard prompt-based paradigms.
More recent work has emphasized the integration of LLMs in graph reasoning [11], noise-robust multimodal pretraining [12], and structured thought modeling [15], expanding the boundaries of explainable and generalizable AI. In computer vision, methods like CPDR [6], DDUNet [7], and causal model-driven super-resolution frameworks such as CausalSR [9] have advanced efficient visual representation learning. Additionally, innovative sensing approaches, such as heartbeat monitoring via RFID and latent diffusion models [14], demonstrate the growing intersection of signal modeling and generative AI.
II Related Work
II-A Bayesian Network Model
Bayesian networks excel at handling uncertainty and predicting rare events [13], both critical challenges in fraud detection. By utilizing probabilistic reasoning, they can evaluate the likelihood of fraud based on available evidence, even when data is incomplete or missing, as illustrated in Fig. 1. Additionally, Bayesian networks are particularly effective in identifying rare fraud cases by incorporating prior knowledge and reasoning under uncertainty, enabling accurate predictions even when the available data is sparse or limited [3].
@@ II-A Bayesian Network Model
II-B Recurrent Neural Networks
@@ III Methods
III-A Dataset
The IBM credit card transaction dataset [2] is a publicly available resource designed for fraud detection research. Each record contains attributes such as transaction amount, card type, and location, along with a label indicating whether the transaction was fraudulent, the details shows at table I. This provides sufficient feature information in each row to train and evaluate models effectively. Notably, the dataset is entirely synthetic and does not correspond to real individuals or financial institutions. It comprises 24 million unique transactions, 6,000 merchants, 100,000 cards, and 30,000 fraudulent cases, accounting for 0.1% of the total transactions. Using this dataset to feed into our GNN model, we would work with approximately 16,000 nodes and 100,000 edges, representing the relationships and interactions within the transaction data.
TABLE I: Dataset Attributes and Types
@@ III-A Dataset
III-B Architecture
@@
In this GNN, each node is characterized by two types of attributes: a concatenated string combining the card number and user name, and the merchant name represented as a string. For the edges, we primarily incorporate the attributes of each transaction as edge features, as shown in Fig. 3. This approach ensures that all relevant information is effectively captured within the graph.
@@
In our model, we use RGCN (Relational Graph Convolutional Network) as the core component. RGCN is a type of GNN specifically designed for handling heterogeneous graphs, which have different types of nodes and edges, as illustrated in Fig. 4. It is capable of learning on multi-relational graphs. In fraud detection scenarios, the data is often heterogeneous. Compared to traditional GCN, RGCN can model different types of relationships more accurately, improving predictive performance.
@@ IV Experiments
The loss curve shown as Fig. 5. The training and testing loss curves in the figure show a sharp decline early on, indicating that the model is learning quickly. Initially, both the training and testing losses are relatively high, but they drop rapidly during the first few epochs. After around epoch 30, the losses stabilize, and both curves approach near-zero values. The training loss (blue curve) decreases more sharply and levels off, suggesting that the model is effectively learning the patterns in the data. The testing loss (orange curve) follows a similar trend but experiences more minor fluctuations, which is typical due to the nature of the testing set. Overall, the steady decrease in loss values across both training and testing phases demonstrates that the model is converging well within the first 200 epochs, even though the testing loss shows slight variability. The final accuracy can achieve 0.9988%, which is highly approach 100%.